OpenStudy (anonymous):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

x + (2x+8)=56

OpenStudy (anonymous):

16 assume one's age to be x um tony's age is x and bel's age is 2x +8

OpenStudy (shane_b):

T+B=56 B=8+2T Using substitution: T + (8+2T)=56 3T=48 T=16 16+B=56 B=40